OpenStudy (anonymous):

What do u guys think the answer is to.. Tell whether the polynomial is completely factored, if not factor it. x2-6x+9=(x-3)2

OpenStudy (anteater):

Is that x^2 - 6x + 9 = (x - 3)^2 ?

OpenStudy (anteater):

If so, then yes. Since (x - 3)^2 = (x - 3)(x-3) = x^2 - 3x - 3x + 9 = x^2 - 6x + 9

OpenStudy (anonymous):

Hello tylermckinney, The problem is not completely factored. The answer to your question is (x−3)^2 Hope this helps!!!!